OpenStudy (anonymous):

Ada gets water from the tap that is Forty-eight degrees Fahrenheit. She puts in on the stove to boil for tea. The graph below shows the relationship between time and temperature of the water over the course of 4 minutes. What is the rate of change for this function?

OpenStudy (smilewithmehun):

Can you show the graph that they're speaking of?

OpenStudy (anonymous):

it wont let me put it

OpenStudy (anonymous):

212 - 48 = 164. It is over 4 minutes so the line has a slope or rate of change of 41°F per minute

OpenStudy (anonymous):

that answer is not in the chocies

OpenStudy (anonymous):

well I don't know what else to do other than subtracting the starting and ending temperatures and then taking that difference and dividing it by the time to get the rate

OpenStudy (smilewithmehun):

it goes up 12 degrees every minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how is it 12 per minute. I took the boiling point of 212 and the starting temp of 48 and subtracted them to get 164. I then divided that by 4 to get 41 per minute so basically what I was doing was this: r = d/t with d being the difference between the starting and ending temperatures

OpenStudy (smilewithmehun):

Look at the graph @caters, for each minute the Farenheit goes up 12 degrees 60-48=12 72-60=12 84-72=12 96-84=12

OpenStudy (clalgee):

Welcome to Openstudy!